# ERDŐS–PÓSA PROPERTY OF CYCLES THAT ARE FAR APART

### v2

@@ 1 Introduction
In 1965, Erdős and Pósa [2] showed that every graph $G$ either contains $k$ vertex-disjoint cycles or contains a set $X$ of $\mathcal{O}(k\log k)$ vertices such that $G-X$ has no cycles. This result has inspired an enormous amount of work on the so-called Erdős-Pósa property of combinatorial objects. For example, using their Grid Minor Theorem, Robertson and Seymour [5] proved the following generalization: For every planar graph $H$, there exists a function $f_{H}(k)$ such that every graph $G$ contains either $k$ vertex-disjoint subgraphs each having an $H$ minor, or a set $X$ of at most $f_{H}(k)$ vertices such that $G-X$ has no $H$ minor.
@@
We prove Theorem 1 for $f(k)\in\mathcal{O}(k^{18}\operatorname{polylog}k)$ and $g(d)=19d$. The statement of Theorem 1 was also posed as a conjecture by Ahn et al. [1], who solved the cases $k=2$ and arbitrary $d$ as well as $d=1$ and arbitrary $k$.
@@ 2 Tools
$\leavevmode\color[rgb]{0.76,0.13,0.28}s(k):=\begin{cases}\lceil 4k(\log k+\log\log k+4)\rceil&\textrm{if $k\geq 2$}\\ 2&\textrm{if $k=1$.}\end{cases}$
@@
There exists a function $\ell^{\star}:\mathbb{N}\times\mathbb{N}\to\mathbb{N}$ such that the following is true. For every $k,c\in\mathbb{N}$ with $k,c\geq 1$, for every forest $F$ and every collection $\mathcal{A}$ of non-null subgraphs of $F$, each having at most $c$ components, either
@@
Let $\ell^{\star}$ be the function given by Theorem 3. Gyárfás and Lehel [4] give a complete proof of Theorem 3 when $F$ is a path and sketch the proof for the general case in which $F$ is a forest. For completeness, we give a proof in the appendix. We show that we can take $\ell^{\star}$ so that $\ell^{\star}(k,3)=\mathcal{O}(k^{18})$, which gives a bound of $f(k)\in\mathcal{O}(k^{18}\operatorname{polylog}k)$ in Theorem 1.
@@ 3 The Proof
Let $G$ be a graph. We use the notations $\leavevmode\color[rgb]{0.76,0.13,0.28}V(G)$ and $\leavevmode\color[rgb]{0.76,0.13,0.28}E(G)$ to denote the vertex set of $G$ and the edge set of $G$, respectively. For a set $S$, $\leavevmode\color[rgb]{0.76,0.13,0.28}G[S]$ denotes the subgraph of $G$ induced by the vertices in $S\cap V(G)$, and $\leavevmode\color[rgb]{0.76,0.13,0.28}G-S:=G[V(G)\setminus S]$. For a set $Z$ of edges of $G$, $G-Z$ denotes the subgraph of $G$ obtained by removing the edges in $Z$ from $G$. For two subsets $A$ and $B$ of $V(G)$, we say that an edge $uv$ of $G$ with $u\in A$ and $v\in B$ is between $A$ and $B$.
The length, $\leavevmode\color[rgb]{0.76,0.13,0.28}\operatorname{len}(P)$, of a path $P$ is the number of edges in $P$. The distance between two vertices $x$ and $y$ of $G$, denoted by $\leavevmode\color[rgb]{0.76,0.13,0.28}\operatorname{dist}_{G}(x,y)$, is the length of a shortest path in $G$ with endpoints $x$ and $y$ if there is one, or $\infty$ if there is none. For nonempty subsets $X$ and $Y$ of $V(G)$, define $\leavevmode\color[rgb]{0.76,0.13,0.28}\operatorname{dist}_{G}(X,Y):=\min\{\operatorname{dist}_{G}(x,y):(x,y)\in X\times Y\}$. For an integer $r\geq 0$ and a vertex $x$ of $G$, the ball of radius $r$ around $x$ is $\leavevmode\color[rgb]{0.76,0.13,0.28}B_{G}(x,r):=\{y\in V(G):\operatorname{dist}_{G}(x,y)\leq r\}$. For a subset $X$ of $V(G)$, $\leavevmode\color[rgb]{0.76,0.13,0.28}B_{G}(X,r):=\bigcup_{x\in X}B_{G}(x,r)$.
A graph $G$ is unicyclic if $G$ contains at most one cycle. Let $r$ be a nonnegative integer and let $G$ be a graph. A cycle $C$ in $G$ is $r$-unicyclic in $G$ if $G[B_{G}(V(C),r)]$ is unicyclic. (Note that every cycle in $G$ is 00-unicyclic.)
@@
    $G$ has a cycle $C^{\prime}$ of length at most $26r+$ 2 with $V(C^{\prime})\subseteq B_{G}(V(C),3r)$.
@@
Let $C_{0}:=C$. We construct inductively a sequence of pairs $(C_{i},Q_{i})_{i\geq 0}$ such that (1) $C_{i}$is a cycle in $G$; (2) $Q_{i}\subseteq C_{i}\cap C$, $Q_{i}$ is connected and contains at least one edge; and (3) $C_{i}$contains at most $14r+$ 1 edges not in $Q_{i}$. Note that the first two conditions imply that there is a path $P_{i}\subseteq C_{i}$ such that $P_{i}$ and $Q_{i}$ are edge-disjoint and $C_{i}=P_{i}\cup Q_{i}$. Thus, $P_{i}$ is a path of length at most $14r+$ 1 with both endpoints in $C$ which implies that all vertices of $P_{i}$ are at distance at most $2r$ from $C$ in $G$. In particular, $V(C_{i})\subseteq B_{G}(V(C),2r)$.
Let $i\geq 0$ and suppose that we already have defined $C_{i}$. If $B_{G}(V(C_{i}),r)$ is unicyclic, then $C_{i}$ witnesses (a), and we are done.
Now suppose that $G[B_{G}(V(C_{i}),r)]$ contains a cycle $D$ distinct from $C_{i}$. If $C_{i}$ has a chord $xy$, then we define a path $P:=xy$. Otherwise, we construct a path $P$ as follows. Since $C_{i}$ has no chord, $D$ contains a vertex not in $C_{i}$. Let $u$ be a vertex of $D$ that is at maximum distance from $C_{i}$ in $G$ among all vertices in $D$. Let $P(u)$ be a shortest path in $G$ from $V(C_{i})$ to $u$. Since $u$ is adjacent to two vertices in $D$ there is a neighbor $v$ of $u$ in $D$ that is not in $P(u)$. Let $P(v)$ be a shortest path from $V(C_{i})$ to $v$ in $G$. Note that both $P(u)$ and $P(v)$ have length at most $r$. By the choice of $u$, the path $P(v)$ does not contain $u$. If $P(u)\cup P(v)\cup\{uv\}$ contains a cycle, say $E$, then $E$ has length at most $12r+$ 1 and $V(E)\subseteq B_{G}(V(C_{i}),r)\subseteq B_{G}(V(C),3r)$, so $E$ witnesses (b). Thus, we assume that $P:=P(u)\cup P(v)\cup\{uv\}$ contains no cycle, so $P$ must be a path.
@@
    Exactly one endpoint of $P$ is in $V(P_{i})$. In this case, $C_{i}\cup P$ has two cycles that each contain $P$. Each edge of $P_{i}$ belongs to exactly one of these two cycles. Therefore one of these cycles uses at most $\lfloor\frac{4r+1}{2}\rfloor=2r$ edges of $P_{i}$. We take $C_{i+1}$ to be this cycle and define $Q_{i+1}=C_{i+1}\cap Q_{i}$.
@@
      for each $C\in\mathcal{C}$ there exists a $C$-rooted BFS-spanning unicycle $U$ of $G[B_{G}(V(C),r)]$ such that $Y$ contains at least one endpoint of each edge in $G[B_{G}(V(C),r)]-E(U)$;
@@
Lemma 5 is an easy consequence of the next two technical lemmas, Lemmas 6 and 7, that address i and ii somewhat separately. Thus, we defer the proof of Lemma 5 until after the proofs of Lemmas 6 and 7.
@@
Let $d,r,k$ be integers such that $1\leq d\leq r$ and $k\geq 1$. Let $G$ be a graph and let $C$ be a $d$-unicyclic cycle in $G$. Let $U$ be a $C$-rooted spanning BFS-unicycle of $G_{r}:=G[B_{G}(V(C),r)]$. Then either:
@@
Consider the auxiliary graph $H$ with vertex set $V(H):=E(G_{r})\setminus E(U)$ in which two distinct elements $e$ and $e^{\prime}$ are adjacent in $H$ if $\operatorname{dist}_{G}(V(P_{e}),V(P_{e^{\prime}}))\leq d$. Let $I$ be a maximal independent set in $H$. We split the proof into two cases: (1) $|I|\geq k+\frac{1}{2}s(k)$, in which case we show that a holds; or (2) $|I|<k+\frac{1}{2}s(k)$, in which case we show that b holds.
@@
By Theorem 2, $G^{\prime}$ contains a set $\mathcal{D}$ of $k$ pairwise vertex-disjoint cycles. We now show that $\mathcal{D}$ is a $d$-packing of cycles in $G$. Let $D$ and $D^{\prime}$ be two distinct cycles in $\mathcal{D}$. Let $v$ and $v^{\prime}$ be vertices of $D$ and $D^{\prime}$, respectively, such that $\operatorname{dist}_{G}(v,v^{\prime})=\operatorname{dist}_{G}(V(D),V(D^{\prime}))$. We must show that $\operatorname{dist}_{G}(v,v^{\prime})>d$.
@@
Since $V(P)\cap V(C)=\{v\}$, $v^{\prime}\notin V(C)$. Therefore $v^{\prime}$ is an internal vertex of $V(P_{e^{\prime}})$ for some $e^{\prime}\in I\setminus J$. Since every internal vertex of $P_{e^{\prime}}$ has degree $2$ in $G^{\prime}$, this implies that $P_{e^{\prime}}\subseteq D^{\prime}$. Since $C$ is $d$-unicyclic in $G$, every path in $G[B_{G}(V(C),d)]$ from $V(C)$ to $v^{\prime}$ contains an endpoint of $P_{e^{\prime}}$. In particular, $P$ contains an endpoint of $P_{e^{\prime}}$. Since $V(P)\cap V(C)=\{v\}$, this implies that $v$ is an endpoint of $P_{e^{\prime}}$. Then $v$ is a vertex of $P_{e^{\prime}}\subseteq D^{\prime}$, which contradicts the fact that $D$ and $D^{\prime}$ are vertex disjoint.
It remains to consider the case when $|I|<k+\frac{1}{2}s(k)$. We now define $X\subseteq V(G)$ and $Y\subseteq V(G)$ and show that these sets satisfy b.
@@
Without loss of generality, suppose $B_{G}(\{u^{\prime},v^{\prime}\},r+d)$ contains a vertex $x$ of $P_{u}$. Then $B_{G}(x,r)$ contains all $U$-descendants of $u$. Therefore, $B_{G}(\{u^{\prime},v^{\prime}\},2r+d)$ contains all $U$-descendants of $u$. We add $u$ to $Y$. Processing each edge $uv\in E(G_{r})\setminus E(U)$ this way produces a set $Y$ so that $X$ and $Y$ satisfy b. ∎
@@
For each $i\in[2]$ and each $v\in V(U_{i})$, let $P_{i,v}$ be the shortest path from $v$ to $V(C_{i})$ in $U_{i}$. Let $Y:=B_{G}(V(C_{1}),r)\cap B_{G}(V(C_{2}),r)$. For each $x$ in $Y$, define $Q_{x}:=P_{1,x}\cup\{x\}\cup P_{2,x}$ and define $P_{x}$ to be a shortest path from $V(C_{1})$ to $V(C_{2})$ in $Q_{x}$.
Let $x$ be an element of $Y$. Since $\operatorname{dist}(V(C_{1}),V(C_{2}))>d\geq 0$, the cycles $C_{1}$ and $C_{2}$ are vertex disjoint. The path $P_{x}$ is a non-trivial path between $V(C_{1})$ and $V(C_{2})$ of length at least $1$ and at most $\operatorname{len}(P_{1,x})+\operatorname{len}(P_{2,x})\leq 2r$. ($P_{x}$ does not necessarily contain the vertex $x$.) We say that $x\in Y$ is uninteresting if $P_{x}$ contains a $U_{i}$-descendant of some vertex in $Y_{i}$ for some $i\in[2]$, otherwise $x$ is interesting.
Let $H$ be an auxiliary graph whose vertex set is the set of interesting elements in $Y$ and in which two distinct vertices $x$ and $x^{\prime}$ are adjacent if and only if $\operatorname{dist}_{G}(V(P_{x}),V(P_{x^{\prime}}))\leq d$. Let $I$ be a maximal independent set in $H$. We split the proof into two cases. If $|I|\geq\frac{1}{2}s(k)$ we will show that outcome a holds. Otherwise, we show that outcome b holds.
First consider the case in which $|I|\geq\tfrac{1}{2}s(k)$. Let $G^{\prime}:=C_{1}\cup C_{2}\cup\bigcup_{x\in I}P_{x}$. For distinct $x,x^{\prime}\in I$, $xx^{\prime}$ is not an edge of $H$, so $\operatorname{dist}_{G}(V(P_{x}),V(P_{x^{\prime}}))>d\geq 1$. This together with the fact that $C_{1}$ is vertex-disjoint from $C_{2}$ gives that $G^{\prime}$ contains only vertices of degree $2$ and degree $3$. Recall that for each $x\in Y$, the path $P_{x}$ contains two different endpoints. Therefore, the degree-$3$ vertices of $G^{\prime}$ are exactly the endpoints of paths in $\{P_{x}:x\in I\}$. This implies that $G^{\prime}$ contains $2|I|\geq s(k)$ vertices of degree $3$. By Theorem 2, $G^{\prime}$ contains a set $\mathcal{D}$ of $k$ pairwise vertex-disjoint cycles.

### v3

@@ 1 Introduction
In 1965, Erdős and Pósa [2] showed that every graph $G$ either contains $k$ vertex-disjoint cycles or contains a set $X$ of $\mathcal{O}(k\log k)$ vertices such that GXG-X has no cycles. This result has inspired an enormous amount of work on the so-called Erdős-Pósa property of combinatorial objects. For example, using their Grid Minor Theorem, Robertson and Seymour [5] proved the following generalization: For every planar graph $H$, there exists a function $f_{H}(k)$ such that every graph $G$ contains either $k$ vertex-disjoint subgraphs each having an $H$ minor, or a set $X$ of at most $f_{H}(k)$ vertices such that GXG-X has no $H$ minor.
@@
We prove Theorem˜1 for $f(k)\in\mathcal{O}(k^{18}\operatorname{polylog}k)$ and $g(d)=19d$. The statement of Theorem˜1 was also posed as a conjecture by Ahn et al. [1], who solved the cases $k=2$ and arbitrary $d$ as well as $d=1$ and arbitrary $k$.
@@ 2 Tools
$\color[rgb]{0.76,0.13,0.28}s(k):=\begin{cases}\lceil 4k(\log k+\log\log k+4)\rceil&\textrm{if $k\geq 2$}\\ 2&\textrm{if $k=1$.}\end{cases}$
@@
There exists a function $\ell^{\star}:\mathbb{N}\times\mathbb{N}\to\mathbb{N}$ such that the following is true. For every $k,c\in\mathbb{N}$ with $k,c\geq 1$, for every forest $F$ and every set $\mathcal{A}$ of non-null subgraphs of $F$, each having at most $c$ components, either
@@
Let $\ell^{\star}$ be the function given by Theorem˜3. Gyárfás and Lehel [4] give a complete proof of Theorem˜3 when $F$ is a path and sketch the proof for the general case in which $F$ is a forest. For completeness, we give a proof in Section˜5. We show that we can take $\ell^{\star}$ so that $\ell^{\star}(k,3)=\mathcal{O}(k^{18})$, which gives a bound of $f(k)\in\mathcal{O}(k^{18}\operatorname{polylog}k)$ in Theorem˜1.
3 Definitions and outline of the proof
In this section we give an overview of the proof of Theorem˜1. We start with basic definitions. Let $G$ be a graph. We use the notations $\color[rgb]{0.76,0.13,0.28}V(G)$ and $\color[rgb]{0.76,0.13,0.28}E(G)$ to denote the vertex set of $G$ and the edge set of $G$, respectively. For a set $S$, $\color[rgb]{0.76,0.13,0.28}G[S]$ denotes the subgraph of $G$ induced by the vertices in $S\cap V(G)$, and $\color[rgb]{0.76,0.13,0.28}G-S:=G[V(G)\setminus S]$. For a set $Z$ of edges of $G$, GZG-Z denotes the subgraph of $G$ obtained by removing the edges in $Z$ from $G$. For two subsets $A$ and $B$ of $V(G)$, we say that an edge $uv$ of $G$ with $u\in A$ and $v\in B$ is between $A$ and $B$.
The length, $\color[rgb]{0.76,0.13,0.28}\operatorname{len}(P)$, of a path $P$ is the number of edges in $P$. The distance between two vertices $x$ and $y$ of $G$, denoted by $\color[rgb]{0.76,0.13,0.28}\operatorname{dist}_{G}(x,y)$, is the length of a shortest path in $G$ with endpoints $x$ and $y$ if there is one, or $\infty$ if there is none. For nonempty subsets $X$ and $Y$ of $V(G)$, define $\color[rgb]{0.76,0.13,0.28}\operatorname{dist}_{G}(X,Y):=\min\{\operatorname{dist}_{G}(x,y):(x,y)\in X\times Y\}$. For an integer $r\geq 0$ and a vertex $x$ of $G$, the ball of radius $r$ around $x$ is $\color[rgb]{0.76,0.13,0.28}B_{G}(x,r):=\{y\in V(G):\operatorname{dist}_{G}(x,y)\leq r\}$. For a subset $X$ of $V(G)$, $\color[rgb]{0.76,0.13,0.28}B_{G}(X,r):=\bigcup_{x\in X}B_{G}(x,r)$.
A graph $G$ is unicyclic if $G$ contains at most one cycle. (In particular, if a unicyclic graph contains a cycle then this cycle must be chordless.) Let $r$ be a nonnegative integer and let $G$ be a graph. A cycle $C$ in $G$ is $r$-unicyclic in $G$ if $G[B_{G}(V(C),r)]$ is unicyclic. (Note that every induced cycle in $G$ is 0-unicyclic.)
We continue with an overview of the proof of Theorem˜1. Let $k\geq 1$, $d\geq 1$, and let $G$ be a graph. Suppose that $G$ does not contain a $d$-packing of $k$ cycles (otherwise, we are done). The goal is to find a small number of balls (bounded in $k$) of small radii (bounded in $d$) that hit all the cycles in $G$.
First, we take a maximal $2d$-packing $\mathcal{C}=\{C_{1},\ldots,C_{p}\}$ of cycles that are $d$-unicyclic in $G$. Since $G$ has no $d$-packing of $k$ cycles, $p<k$. Next, in a simple iterative process, we build a $d$-packing $\mathcal{D}=\{D_{1},\ldots,D_{q}\}$ of ’short’ cycles in $G$, of length at most $6d+2$. Again, we have $q<k$. The key property of these two collections is that every cycle in $G$ contains a vertex in $B_{G}(\bigcup_{i\in[p]}V(C_{i}),4d)\cup B_{G}(\bigcup_{i\in[q]}V(D_{i}),4d)$. This is a consequence of Lemma˜4, which states that for every cycle $C$ in $G$, either the ball of radius $2d$ around $C$ contains a $d$-unicyclic cycle or the ball of radius $3d$ around $C$ contains a short cycle.
We choose one vertex from each cycle in $\mathcal{D}$ and let $X_{0}$ be the set of chosen vertices. Note that since all cycles in $\mathcal{D}$ are short, we have $B_{G}(\bigcup_{i\in[q]}V(D_{i}))\subseteq B_{G}(X_{0},7d+1)$.
Next, we aim to hit all the cycles in $G$ contained in $B_{G}(V(C_{i}),6d)$, for each $i\in[p]$. Fix $i\in[p]$ and fix a $C_{i}$-rooted spanning BFS-unicycle $U_{i}$ of $G_{i}:=G[B_{G}(V(C_{i}),6d)]$ (see next section for the definition). Each edge $e$ in $E(G_{i})\setminus E(U_{i})$ together with two paths in $U_{i}$ from the endpoints of $e$ to $V(C_{i})$ induces a subgraph $P_{e}$ of $G$ that is a path or a ’lollipop’ in $G_{i}$. One can prove that either there are many such subgraphs $P_{e}$ that are $d$-apart in $G$ or there is a subset of vertices $Z_{i}$ such that $B_{G}(Z_{i},13d)$ hits every edge $e$ in $E(G_{i})\setminus E(U_{i})$. In the former case, this subfamily of paths $\{P_{e}\}$ together with cycles in $\mathcal{C}$ form a ’frame’, a subgraph of $G$ where all vertices have degree $2$ or $3$, and with many vertices of degree $3$. One can then apply Simonovits’s Theorem, see Theorem˜2, which gives a family of $k$ vertex-disjoint cycles in the frame. This packing of cycles in the frame translates into a $d$-packing of $k$ cycles in $G$, a contradiction. Thus, we obtain the second outcome, that is, a subset $Z_{i}$ of vertices in $G$ such that $B_{G}(Z_{i},13d)$ contains an endpoint of every edge $e$ in $E(G_{i})\setminus E(U)$. This argument is encapsulated in Lemma˜6. We then let $X_{1}:=\bigcup_{i\in[p]}Z_{i}$.
With a similar argument, which involves another application of Simonovits Theorem, and which is enclosed in Lemma˜7, we construct a set $X_{2}\subseteq V(G)$ of bounded size such that all vertices of $G$ that lie in the intersection of at least two sets from $\{B_{G}(V(C_{i}),6d):i\in[p]\}$ are contained in $B_{G}(X_{2},13d)$.
At this point, we are in the following situation: We defined bounded-size sets $X_{0},X_{1},X_{2}$, and we let $\widehat{Y}$ be the union of the balls centered on the vertices in these sets (with the appropriate bounded radius). After the removal of $\widehat{Y}$, the graph $G$ has no cycle that is at distance more than $4d$ from all cycles in $\{C_{1},\ldots,C_{p}\}$. Thus, the graph
$F_{0}:=\textstyle{G-\left(\widehat{Y}\cup\bigcup_{i\in[p]}B_{G}(V(C_{i}),4d)\right)}$
is a forest. Similarly, for each $i\in[p]$, we made sure that every cycle in the $6d$-ball around $C_{i}$ is hit by $\widehat{Y}$, thus
$F_{i}:=G[B_{G}(V(C_{i}),6d)]-\widehat{Y}$
is a forest as well. Now, every cycle in $G$ that avoids $\widehat{Y}$ will necessarily intersect at least two of the forests $F_{0},F_{1},\dots,F_{p}$ (possibly many of them). Moreover, every such cycle $C$ must intersect $F_{0}$ (which we think of as the "forest outside"), thanks to the properties of $\widehat{Y}$, namely, $\widehat{Y}$ contains all vertices of $G$ that are contained in the intersection of two $6d$-balls around two cycles in $\{C_{1},\ldots,C_{p}\}$. In other words, if our cycle $C$ enters the $6d$-ball around some cycle $C_{i}$, then it has to exit it via $F_{0}$. This naturally leads us to consider the ’outside segments’ of $C$, which are (roughly) the parts of $C$ contained in $F_{0}$, and the ’inside segments’ of $C$ that belong to some $F_{i}$ with $i\in[p]$.
The final step of the proof, which is perhaps the most technical one, is to consider the $d$-balls around all the segments defined by all cycles of $G$ that avoid $\widehat{Y}$, and use them in an application of Theorem˜3—a powerful generalization of the Helly property of subtrees of a tree—using crucially that the number of ’dimensions’ in our application is bounded: one for each of the forests $F_{0},F_{1},\dots,F_{p}$. The latter result either gives a bounded number of vertices hitting the $d$-balls of all these segments, which we use as centers of balls to hit the remaining cycles in $G$. Or it gives a large number of segments with pairwise disjoint $d$-balls. In the latter case, we use these to find $k$ cycles pairwise at distance more than $d$ using an application of Simonovits Theorem again.
We remark that the second part of the above outline, once $\widehat{Y}$ has been built, is a simplified view of the actual proof that does not mention various subtleties but it is good enough to get a first intuitive picture of the proof.
@@ 4 The Proof
    $G$ has a cycle $C^{\prime}$ of length at most $6r+2$ with $V(C^{\prime})\subseteq B_{G}(V(C),3r)$.
@@
Let $C_{0}:=C$. We construct inductively a sequence of pairs $(C_{i},Q_{i})_{i\geq 0}$ such that (1) $C_{i}$is a cycle in $G$; (2) $Q_{i}\subseteq C_{i}\cap C$, $Q_{i}$ is a path and contains at least one edge; and (3) $C_{i}$contains at most $4r+1$ edges not in $Q_{i}$. Note that the first two conditions imply that there is a path $P_{i}\subseteq C_{i}$ such that $P_{i}$ and $Q_{i}$ are edge-disjoint and $C_{i}=P_{i}\cup Q_{i}$. Thus, $P_{i}$ is a path of length at most $4r+1$ with both endpoints in $C$ which implies that all vertices of $P_{i}$ are at distance at most $2r$ from $C$ in $G$. In particular, $V(C_{i})\subseteq B_{G}(V(C),2r)$.
Let $i\geq 0$ and suppose that we already have defined $(C_{i},Q_{i})$. If $B_{G}(V(C_{i}),r)$ is unicyclic, then $C_{i}$ witnesses (a), and we are done.
Now suppose that $G[B_{G}(V(C_{i}),r)]$ contains a cycle $D$ distinct from $C_{i}$. If $C_{i}$ has a chord $xy$, then we define a path $P:=xy$. Otherwise, we construct a path $P$ as follows. Since $C_{i}$ has no chord, $D$ contains a vertex not in $C_{i}$. Let $u$ be a vertex of $D$ that is at maximum distance from $C_{i}$ in $G$ among all vertices in $D$. Let $P(u)$ be a shortest path in $G$ from $V(C_{i})$ to $u$. Since $u$ is adjacent to two vertices in $D$ there is a neighbor $v$ of $u$ in $D$ that is not in $P(u)$. Let $P(v)$ be a shortest path from $V(C_{i})$ to $v$ in $G$. Note that both $P(u)$ and $P(v)$ have length at most $r$. By the choice of $u$, the path $P(v)$ does not contain $u$. See Figure 1. If $P(u)\cup P(v)\cup\{uv\}$ contains a cycle, say $E$, then $E$ has length at most $2r+1$ and $V(E)\subseteq B_{G}(V(C_{i}),r)\subseteq B_{G}(V(C),3r)$, so $E$ witnesses (b). Thus, we assume that $P:=P(u)\cup P(v)\cup\{uv\}$ contains no cycle, so $P$ must be a path.
Figure 1: Proof of Lemma˜4: on the left $P(u)\cup P(v)\cup\{uv\}$ contains a cycle; on the right $P(u)\cup P(v)\cup\{uv\}$ is a path with both ends in $V(C_{i})$.
@@
    Exactly one endpoint of $P$ is in $V(P_{i})$. In this case, $C_{i}\cup P$ has two cycles that each contain $P$. Each edge of $P_{i}$ belongs to exactly one of these two cycles. Therefore one of these cycles uses at most $\lfloor\frac{4r+1}{2}\rfloor=2r$ edges of $P_{i}$. We take $C_{i+1}$ to be this cycle and define $Q_{i+1}=C_{i+1}\cap Q_{i}$. Note that $\operatorname{len}(Q_{i+1})<\operatorname{len}(Q_{i})$ because $C_{i+1}$ does not contain one of the endpoints of $Q_{i}$.
@@
      for each $C\in\mathcal{C}$ there exists a $C$-rooted spanning BFS-unicycle $U$ of $G[B_{G}(V(C),r)]$ such that $Y$ contains at least one endpoint of each edge in $G[B_{G}(V(C),r)]-E(U)$;
@@
Lemma˜5 is an easy consequence of the next two technical lemmas, Lemmas˜6 and 7, that address ˜i and ˜ii somewhat separately. Thus, we defer the proof of Lemma˜5 until after the proofs of Lemmas˜6 and 7.
@@
Let $d,r,k$ be integers such that $r\geq d\geq 1$ and $k\geq 1$. Let $G$ be a graph and let $C$ be a $d$-unicyclic cycle in $G$. Let $U$ be a $C$-rooted spanning BFS-unicycle of $G_{r}:=G[B_{G}(V(C),r)]$. Then either:
@@
Figure 2: Illustration of outcome ˜b of Lemma˜6: $C$ is a $d$-unicyclic cycle in $G$. $U$ is a $C$-rooted spanning BFS-unicycle of $G_{r}=G[B_{G}(V(C),r)]$. The edges of $E(G_{r})\setminus E(U)$ are highlighted, and one of endpoint of each such edge is highlighted. The set of these highlighted endpoints is $Y$. Finally, $X$ is the set of centers of a small number of balls (four in this figure) whose union contains all $U$-descendants of vertices in $Y$.
@@
Consider the auxiliary graph $H$ with vertex set $V(H):=E(G_{r})\setminus E(U)$ in which two distinct elements $e$ and $e^{\prime}$ are adjacent in $H$ if $\operatorname{dist}_{G}(V(P_{e}),V(P_{e^{\prime}}))\leq d$. Let $I$ be a maximal independent set in $H$. We split the proof into two cases: (1) $|I|\geq k+\frac{1}{2}s(k)$, in which case we show that ˜a holds; or (2) $|I|<k+\frac{1}{2}s(k)$, in which case we show that ˜b holds.
@@
By Theorem˜2, $G^{\prime}$ contains a set $\mathcal{D}$ of $k$ pairwise vertex-disjoint cycles. We now show that $\mathcal{D}$ is a $d$-packing of cycles in $G$. Let $D$ and $D^{\prime}$ be two distinct cycles in $\mathcal{D}$. Let $v$ and $v^{\prime}$ be vertices of $D$ and $D^{\prime}$, respectively, such that $\operatorname{dist}_{G}(v,v^{\prime})=\operatorname{dist}_{G}(V(D),V(D^{\prime}))$. We must show that $\operatorname{dist}_{G}(v,v^{\prime})>d$.
@@
Since $V(P)\cap V(C)=\{v\}$, $v^{\prime}\notin V(C)$. Therefore $v^{\prime}$ is an internal vertex of $V(P_{e^{\prime}})$ for some $e^{\prime}\in I\setminus J$. Since every internal vertex of $P_{e^{\prime}}$ has degree $2$ in $G^{\prime}$, this implies that $P_{e^{\prime}}\subseteq D^{\prime}$. Since $C$ is $d$-unicyclic in $G$, every path in $G[B_{G}(V(C),d)]$ from $V(C)$ to $v^{\prime}$ contains an endpoint of $P_{e^{\prime}}$. In particular, $P$ contains an endpoint of $P_{e^{\prime}}$. Since $V(P)\cap V(C)=\{v\}$, this implies that $v$ is an endpoint of $P_{e^{\prime}}$. Then $v$ is a vertex of $P_{e^{\prime}}\subseteq D^{\prime}$, which contradicts the fact that $D$ and $D^{\prime}$ are vertex-disjoint.
It remains to consider the case when $|I|<k+\frac{1}{2}s(k)$. We now define $X\subseteq V(G)$ and $Y\subseteq V(G)$ and show that these sets satisfy ˜b.
@@
Without loss of generality, suppose $B_{G}(\{u^{\prime},v^{\prime}\},r+d)$ contains a vertex $x$ of $P_{u}$. Then $B_{G}(x,r)$ contains all $U$-descendants of $u$. Therefore, $B_{G}(\{u^{\prime},v^{\prime}\},2r+d)$ contains all $U$-descendants of $u$. We add $u$ to $Y$. Processing each edge $uv\in E(G_{r})\setminus E(U)$ this way produces a set $Y$ so that $X$ and $Y$ satisfy ˜b. ∎
@@
For each $i\in[2]$ and each $v\in V(U_{i})$, let $P_{i,v}$ be the shortest path from $v$ to $V(C_{i})$ in $U_{i}$. Let $Y:=B_{G}(V(C_{1}),r)\cap B_{G}(V(C_{2}),r)$. For each $x$ in $Y$, define $Q_{x}:=P_{1,x}\cup P_{2,x}$ and define $P_{x}$ to be a shortest path from $V(C_{1})$ to $V(C_{2})$ in $Q_{x}$.
Let $x$ be an element of $Y$. Since $\operatorname{dist}(V(C_{1}),V(C_{2}))>d\geq 0$, the cycles $C_{1}$ and $C_{2}$ are vertex-disjoint. The path $P_{x}$ is a non-trivial path between $V(C_{1})$ and $V(C_{2})$ of length at least $1$ and at most $\operatorname{len}(P_{1,x})+\operatorname{len}(P_{2,x})\leq 2r$. ($P_{x}$ does not necessarily contain the vertex $x$.) We say that $x\in Y$ is uninteresting if $P_{x}$ contains a $U_{i}$-descendant of some vertex in $Y_{i}$ for some $i\in[2]$, otherwise $x$ is interesting.
Let $H$ be an auxiliary graph whose vertex set is the set of interesting elements in $Y$ and in which two distinct vertices $x$ and $x^{\prime}$ are adjacent if and only if $\operatorname{dist}_{G}(V(P_{x}),V(P_{x^{\prime}}))\leq d$. Let $I$ be a maximal independent set in $H$. We split the proof into two cases. If $|I|\geq\frac{1}{2}s(k)$ we will show that outcome ˜a holds. Otherwise, we show that outcome ˜b holds.
First consider the case in which $|I|\geq\tfrac{1}{2}s(k)$. Let $G^{\prime}:=C_{1}\cup C_{2}\cup\bigcup_{x\in I}P_{x}$. For distinct $^{\prime}\in I, xxxx^{\prime}$ is not an edge of $H$, so $\operatorname{dist}_{G}(V(P_{x}),V(P_{x^{\prime}}))>d\geq 1$. This together with the fact that $C_{1}$ is vertex-disjoint from $C_{2}$ gives that $G^{\prime}$ contains only vertices of degree $2$ and degree $3$. Recall that for each $x\in Y$, the path $P_{x}$ contains two different endpoints. Therefore, the degree-$3$ vertices of $G^{\prime}$ are exactly the endpoints of paths in $\{P_{x}:x\in I\}$. This implies that $G^{\prime}$ contains $2|I|\geq s(k)$ vertices of degree $3$. By Theorem˜2, $G^{\prime}$ contains a set $\mathcal{D}$ of $k$ pairwise vertex-disjoint cycles.